OpenStudy (anonymous):

does anybody know what this means? 3! = 24 9 6 None of the above

OpenStudy (anonymous):

6

OpenStudy (eyust707):

yes that is a factorial symbol

OpenStudy (eyust707):

it means: 3*2*1

OpenStudy (anonymous):

That's how it's written in my mathbook. What does that mean?

OpenStudy (precal):

3 factorial 3x2x1=6

OpenStudy (anonymous):

it means, that you multiply those number so, 3*2=6 and 6*1 = 6

OpenStudy (anonymous):

Oh I thought I was seeing things.. Ha ha. Thanks 4 the help guys :)

OpenStudy (anonymous):

3 factorial would be written as 3 x 2 x 1 Other example: 4! would be 4 x 3 x 2 x 1 = 24

OpenStudy (eyust707):

5! = 5 * 4 * 3 *2 *1 6! = 6 * 5 * 4 * 3 *2 *1 7! = 7 * 6 *5 * 4 * 3 *2 *1

OpenStudy (precal):

5! 5x4x3x2x1=120